# Generalizable and Relightable Gaussian Splatting for Human Novel View Synthesis

### v1

@@ Abstract
We propose GRGS, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under diverse lighting conditions. Unlike existing methods that rely on per-character optimization or ignore physical constraints, GRGS adopts a feed-forward, fully supervised strategy that projects geometry, material, and illumination cues from multi-view 2D observations into 3D Gaussian representations. Specifically, to reconstruct lighting-invariant geometry, we introduce a Lighting-aware Geometry Refinement (LGR) module trained on synthetically relit data to predict accurate depth and surface normals. Based on the high-quality geometry, a Physically Grounded Neural Rendering (PGNR) module is further proposed to integrate neural prediction with physics-based shading, supporting editable relighting with shadows and indirect illumination. Besides, we design a 2D-to-3D projection training scheme that leverages differentiable supervision from ambient occlusion, direct, and indirect lighting maps, which alleviates the computational cost of explicit ray tracing. Extensive experiments demonstrate that GRGS achieves superior visual quality, geometric consistency, and generalization across characters and lighting conditions. Project webpage: https://sypj-98.github.io/grgs/.
@@ 1 Introduction
Human novel view synthesis (NVS) aims to produce photorealistic images of human performers under a specific targeting novel viewpoint, which has a wide range of applications such as immersive telepresence, cinematic production, and AR/VR. To further improve realism and fidelity, relightable rendering, which enables editing lighting during synthesizing novel views, has become an important improvement over traditional NVS, as lighting significantly affects surface appearance, immersive shading, and spatial consistency. Although significant efforts have been devoted to developing effective relightable rendering, it is still a challenging task due to complex light transport in 3D space.
@@
Prior methods [9, 24, 15, 13] typically rely on mesh-based geometry and texture reconstruction to support plausible viewpoint transitions, but they often involve complex reconstruction pipelines and struggle with generating accurate and smooth mesh surfaces, limiting their effectiveness for realistic relighting without extensive post-processing. Recent advances [30, 35, 21] in neural representations have substantially propelled the field of NVS. Specifically, NeRF-based approaches [41, 62, 59, 18, 56, 58, 25, 8, 28, 55] incorporate physically-based rendering (PBR) principles into volumetric radiance fields or signed distance fields (SDFs) for implicit geometry and material encoding, enabling the synthesis of photorealistic images under novel viewpoints and lighting conditions. However, these methods typically require time-consuming per-character optimization and incur high computational costs, resulting in limited rendering speeds that hinder their practicality in real-world applications. Recently, 3D Gaussian Splatting (3DGS) [21] has emerged as an efficient and explicit neural representation, offering an efficient training and inference process while maintaining high visual fidelity. To adapt 3DGS for relighting tasks, researchers [12, 17, 27, 51, 26] augment each 3D Gaussian point with intrinsic properties (geometry and appearance) and employ inverse rendering techniques to estimate light transport. Unfortunately, it remains reliant on iterative inverse rendering pipelines that are inherently ill-posed, resulting in inaccurate estimation of geometry and materials, which degrade the overall rendering quality. Moreover, per-character optimization still prevents them from being applied to relighting applications that require generalization across different characters.
@@
In contrast to 3D methods, 2D image-based relighting approaches employ encoder-decoder architectures (e.g., U-Net) [20, 64, 36, 16, 33, 22] or diffusion models [38, 39, 60] to learn lighting priors from large-scale relighting datasets, producing high-quality relit images under novel illumination in a generalized manner. Despite their efficiency and strong generalization capabilities, the lack of 3D consistent constraints makes 2D image-based methods typically suffer from flickering artifacts when rendering from precisely user-controlled viewpoints. In addition, although these methods often produce visually impressive relighting results, they often ignore physical interpretability due to their reliance on neural networks to implicitly learn physical constraints.
@@
To address these challenges, we propose GRGS, as shown in Fig. 1, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under various lighting through integrating multiple intrinsic attribute priors into 3D Gaussian representations. Unlike existing 3DGS-based methods that apply person-specific optimization, the core idea of GRGS is to adopt a supervised, data-driven strategy that learns to project geometry, material, and illumination cues from multi-view 2D observations onto 3D Gaussian attributes in a feed-forward manner for robust generalization and realistic relighting. Our framework starts by presenting a Lighting-aware Geometry Refinement (LGR) module, which first estimates coarse depth using a stereo-based method and then refines per-Gaussian surface normals to capture smooth and fine-grained details. To obtain reliable geometry under challenging illumination, we synthesize large-scale relit multi-view data, enabling LGR to produce lighting-invariant geometry that mitigates feature mismatches and improves geometric accuracy. The refined geometric priors not only support realistic relighting but also serve as a critical bridge between the 2D image space and the 3D Gaussian domain, facilitating accurate positioning of Gaussians. With the geometry established, we design a Physically Grounded Neural Rendering (PGNR) module that integrates physics-based rendering with neural rendering, which synthesizes realistic shading phenomena, including shadows and indirect illumination, while enforcing physically grounded lighting consistency. PGNR employs geometry-aware decoders to infer Gaussian parameters and intrinsic attributes from the refined geometry in a feed-forward manner. In parallel, a lightweight encoder-decoder processes the high-resolution environment map to estimate direct illumination scaling factors and spherical harmonic coefficients for modeling indirect lighting. By incorporating a physically-based rendering, our framework ensures physically consistent light transport, achieving high-quality and photorealistic relighting results. Besides, the 2D-to-3D project strategy alleviates the computational cost of explicit ray tracing for visibility and global illumination and ensures high efficiency during inference.
In summary, our method makes the following key contributions:
@@
  • We present a Lighting-aware Geometry Refinement module trained on synthetically relit multi-view data to estimate lighting-invariant depth and surface normals, effectively mitigating geometry errors caused by uneven illumination.
  • We design a Physically Grounded Neural Rendering module that combines physics-based rendering with neural rendering, supporting high-quality shading phenomena synthesis while avoiding the computational overhead of explicit ray tracing.
@@ 2 Related Work
Traditional methods [10, 14, 11, 49, 9, 50, 6, 13] propose to sample the reflectance field of a human performer through a LightStage setup comprising controlled lighting systems and dense camera arrays, which can generate photorealistic renderings under novel lighting environments. However, the costly setup and complex person-specific relighting process constrain their widespread applications. With the rapid progress in neural implicit representations [35, 1, 2, 31, 45, 53, 52], neural inverse rendering [3, 41, 62, 59, 4, 18, 56, 58] has emerged as a promising approach for person-specific human relighting, enabling the joint recovery of geometry, material, and illumination from multi-view images captured under arbitrary lighting conditions. Recent variants [8, 43, 54, 28, 26, 7, 47] extend inverse rendering to dynamic performers by integrating a parametric human body template into the learning of implicit fields, enabling temporally coherent reconstruction under varying poses and motions. However, these methods typically require time-consuming person-specific optimization and incur high computational costs, resulting in limited training and rendering speeds. With the advent of explicit point-based 3DGS [21] representation, recent 3DGS-based inverse rendering methods [12, 17, 27, 51] explore the disentanglement of material, geometry, and lighting through the use of Gaussian points, enabling efficient training and fast inference. ARGS [26] pioneers the integration of animatable 3D Gaussians and inverse rendering, delivering visually compelling and relightable full-body human avatars. RGCA [40] leverages dynamic relighting data captured in a LightStage and an explicit mesh template to achieve high-quality relightable head avatars. However, such methods typically rely on time-consuming person-specific optimization. Moreover, performing inverse rendering from multi-view images under arbitrary illumination is inherently ill-posed, making it particularly difficult to accurately disentangle geometry, material, and lighting, ultimately limiting the fidelity of the final relighting results.
@@
With the advancement of deep learning, data-driven methods [36, 42, 64, 48, 33] have achieved impressive results in portrait relighting from a single image, primarily by leveraging convolutional neural networks trained on synthetic data derived from OLAT datasets. However, these methods typically overlook underlying human geometry, leading to physically implausible light–shadow interactions. Recent efforts [34, 16] attempt to address this limitation by estimating human geometry from a single image, but monocular predictions remain insufficiently accurate for reliable relighting. Other approaches [20, 23, 44] attempt to learn light transport directly from 2D shading images. However, Modeling complex 3D light transport in 2D space remains fundamentally challenging, such methods are generally restricted to approximating low-frequency lighting effects and often fail to generalize to unseen illumination conditions. SwitchLight [22] circumvents the need for explicit 3D geometry by introducing physics-based rendering in the image domain, achieving high-quality relighting results. Nevertheless, the absence of an explicit 3D representation limits its ability to handle occlusions and maintain spatial consistency. More recently, diffusion-based generative models [38, 39, 60] have demonstrated impressive relighting performance by leveraging strong pre-trained priors and high-quality lighting datasets, producing visually striking results with high generalization ability. However, both the learned priors from pre-training and the lighting supervision often lack physically grounded constraints, limiting the physical interpretability and realism of the generated outputs.
@@ 3 Method
As illustrated in Fig. 2, our proposed GRGS achieves generalizable and relightable human novel view synthesis via two core modules: (1) Lighting-aware Geometry Refinement (LGR) module and (2) Physically Grounded Neural Rendering (PGNR) module. The core idea is to adopt a supervised, data-driven strategy that projects intrinsic attributes from multi-view 2D observations onto 3D Gaussian representations in a feed-forward manner. Specifically, given a set of sparse-view human images, GRGS first constructs the LGR module to estimate depth using a stereo-based method and then refines per-Gaussian surface normals via a depth-aware refinement, which allows GRGS to mitigate geometry errors caused by uneven illumination (Sec. 3.1). Next, building on this lighting-invariant geometry, the PGNR module fuses physics-based rendering with neural rendering to synthesize realistic relighting phenomena. It leverages geometry-aware decoders and environment map encoding to predict direct illumination scaling factors and spherical harmonic coefficients, avoiding the computational cost of explicit ray tracing (Sec. 3.2). To ensure strong generalization and realistic rendering, we introduce a 2D-to-3D projection training strategy that exploits the diversity of multi-view 2D observations while enforcing geometric consistency in the 3D Gaussian space (Sec. 3.3).
3.1 Lighting-aware Geometry Refinement
To enable a generalizable and relightable 3D Gaussian framework, fast and accurate geometry reconstruction of the target performer is a critical foundation. This geometry not only determines the center of each Gaussian point but also serves as an essential component for the subsequent relighting stage. However, existing methods [12, 17, 27, 51] typically require tens of minutes of per-scene training to optimize Gaussian centers and still struggle to guarantee high-quality and consistent geometry under various lighting. To address this challenge, we design a Lighting-aware Geometry Refinement module, consisting of stereo-based depth estimation and depth-guided normal refinement, which jointly produce lighting-invariant geometry robust to illumination changes.
@@ 3.1 Lighting-aware Geometry Refinement
Inspired by the generalization strategy of GPS-Gaussian [63], we leverage RAFT-Stereo [29], a stereo-based depth estimator that uses disparity as a geometric constraint across views, enabling consistent depth prediction and improved generalization in diverse human-centric scenarios. However, pretrained depth estimators tend to be sensitive to illumination variations, leading to unreliable geometry estimation and degraded relighting performance. Therefore, we construct a lighting-aware image encoder trained by a large-scale multi-view relit dataset comprising hundreds of high-quality human scans, enabling the network to learn robust reflectance features for accurate depth estimation across diverse lighting conditions. Specifically, given $N$ sparse-view images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ ($\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}$), captured under an arbitrary lighting of a human-centered scene, we select the two source views nearest to the target camera pose, denoted as $\{\hat{\mathbf{I}},\bar{\mathbf{I}}\}$, as input for stereo rectification [37]. The rectified image pair is then passed through a shared lighting-aware feature extractor $\xi_{\text{img}}$ to obtain multi-scale radiance features $\{\hat{\mathbf{F}}_{s},\bar{\mathbf{F}}_{s}\}_{s=1}^{K}$, where $\hat{\mathbf{F}}_{s},\bar{\mathbf{F}}_{s}\in\mathbb{R}^{\frac{H}{2^{s}}\times\frac{W}{2^{s}}\times C}$ denote the feature representations at the $s$-th scale corresponding to the two input images $\{\hat{\mathbf{I}},\bar{\mathbf{I}}\}$, respectively, $K$ is the total number of scales. The radiance features mitigate lighting-induced feature mismatches to facilitate accurate depth estimation, as demonstrated in Sec. 4.2, while guiding the inference of geometry, material, and light transport within the 3D Gaussian representation (Sec. 3.2).
@@
$\mathbf{N}_{c}(u,v)=\frac{\left(\mathbf{X}(u+1,v)-\mathbf{X}(u,v)\right)\times\left(\mathbf{X}(u,v+1)-\mathbf{X}(u,v)\right)}{\parallel\left(\mathbf{X}(u+1,v)-\mathbf{X}(u,v)\right)\times\left(\mathbf{X}(u,v+1)-\mathbf{X}(u,v)\right)\parallel}$ (4)
where $\times$ denotes the cross product. Due to the convex upsampling, the resulting depth maps lack high-frequency geometric details and exhibit grid-like artifacts, as illustrated in Sec. 4.2. To address this, we introduce a normal refinement module $\mathcal{N}$ based on a lightweight U-Net architecture, which leverages multi-scale semantic features $\{\mathbf{F}_{s}\}_{s=1}^{K}$ and coarse normals $\mathbf{N}_{c}$ to predict normal offsets $\Delta\mathbf{N}\in\mathbb{R}^{3}$ and recover fine-grained surface details. The refinement process can be defined as

### v2

@@ Abstract
We propose GRGS, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under diverse lighting conditions. Unlike existing methods that rely on per-character optimization or ignore physical constraints, GRGS adopts a feed-forward, fully supervised strategy projecting geometry, material, and illumination cues from multi-view 2D observations into 3D Gaussian representations. To recover accurate geometry under diverse lighting conditions, we introduce a Lighting-robust Geometry Refinement (LGR) module trained on synthetically relit data to predict precise depth and surface normals. Based on the high-quality geometry, a Physically Grounded Neural Rendering (PGNR) module is further proposed to integrate neural prediction with physics-based shading, supporting editable relighting with shadows and indirect illumination. Moreover, we design a 2D-to-3D projection training scheme leveraging differentiable supervision from ambient occlusion, direct, and indirect lighting maps, alleviating the computational cost of ray tracing. Extensive experiments demonstrate that GRGS achieves superior visual quality, geometric consistency, and generalization across characters and lighting conditions. The project webpage is available at https://sypj-98.github.io/grgs/.
@@ 1 Introduction
Human novel view synthesis (NVS) aims to produce photorealistic images of human performers under a specific targeting novel viewpoint, which has a wide range of applications such as immersive telepresence, cinematic production, and AR/VR. To further improve realism and fidelity, relightable rendering, which enables editing lighting during synthesizing novel views, has become an important improvement over traditional NVS, as lighting significantly affects surface appearance and immersive shading. Although significant efforts have been devoted to developing effective relightable rendering, it is still a challenging task due to complex light transport in 3D space.
@@
Prior methods [debevec2012light, li2013capturing, imber2014intrinsic, guo2019relightables] typically rely on mesh-based geometry and texture reconstruction to support plausible viewpoint transitions, but they often involve complex reconstruction pipelines and struggle with generating accurate and smooth mesh surfaces, limiting their effectiveness for realistic relighting without extensive post-processing. Recent advances [Lombardi:2019, mildenhall2021nerf, kerbl20233d] in neural representations have substantially propelled the field of NVS. Specifically, NeRF-based approaches [srinivasan2021nerv, zhang2021nerfactor, jin2023tensoir, yao2022neilf, li2024tensosdf, chen2022relighting4d, xu2024relightable] incorporate physically-based rendering (PBR) principles into volumetric radiance fields or signed distance fields (SDFs) for implicit geometry and material encoding, enabling the synthesis of photorealistic images under novel viewpoints and lighting conditions. However, these methods typically require time-consuming per-character optimization and incur high computational costs, resulting in limited rendering speeds that hinder their practicality in real-world applications. Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as an efficient and explicit neural representation, offering a fast training and inference process while maintaining high visual fidelity. To adapt 3DGS for relighting tasks, researchers [gao2024relightable, jiang2024gaussianshader, liang2024gs, wu2025deferredgs, li2024animatable] augment each 3D Gaussian point with intrinsic properties (geometry and appearance) and employ inverse rendering techniques to estimate light transport. Unfortunately, it remains reliant on iterative inverse rendering pipelines that are inherently ill-posed, resulting in inaccurate estimation of geometry and materials, which degrades the overall rendering quality. Moreover, per-character optimization still prevents them from being applied to relighting applications that require generalization across different characters.
@@
In contrast to 3D methods, 2D image-based relighting approaches employ encoder-decoder architectures (e.g., U-Net) [pandey2021total, ji2022geometry, mei2023lightpainter, kim2024switchlight] or diffusion models [ponglertnapakorn2023difareli, ren2024relightful, zhang2025scaling] to learn lighting priors from large-scale relighting datasets, producing high-quality relit images under novel illumination in a generalized manner. Despite their efficiency and strong generalization capabilities, the lack of 3D consistent constraints makes 2D image-based methods suffer from flickering artifacts when rendering from precisely user-controlled viewpoints. In addition, although these methods often produce visually impressive relighting results, they often overlook physical interpretability as they entirely rely on neural networks to implicitly capture physical constraints.
@@
To address these challenges, we propose GRGS, as shown in Fig. 1, a generalizable and relightable 3D Gaussian framework for high-fidelity human novel view synthesis under various lighting conditions by integrating multiple intrinsic attribute priors into 3D Gaussian representations. Unlike existing 3DGS-based methods that apply person-specific optimization, the core idea of GRGS is to adopt a supervised, data-driven strategy that learns to project geometry, material, and illumination cues from multi-view 2D observations onto 3D Gaussian attributes in a feed-forward manner for robust generalization and realistic relighting. Our framework starts by presenting a Lighting-robust Geometry Refinement (LGR) module, which first estimates coarse depth using a stereo-based method and then refines per-Gaussian surface normals to capture smooth and fine-grained details. To obtain reliable geometry under challenging illumination, we synthesize large-scale relit multi-view data, enabling LGR to produce geometry consistent across diverse lighting conditions, thereby mitigating feature mismatches and improving geometric accuracy. The refined geometric priors not only support realistic relighting but also serve as a critical bridge between the 2D image space and the 3D Gaussian domain, facilitating accurate positioning of Gaussians. With the geometry established, we design a Physically Grounded Neural Rendering (PGNR) module that integrates physics-based rendering with neural rendering, which synthesizes realistic shading phenomena, including shadows and indirect lighting, while enforcing physically grounded lighting consistency. PGNR employs geometry-aware decoders to infer Gaussian parameters and intrinsic attributes from the refined geometry in a feed-forward manner. In parallel, a lightweight encoder-decoder processes the high-resolution environment map to estimate direct illumination scaling factors and spherical harmonic coefficients for modeling indirect lighting. By incorporating a physically-based rendering, our framework ensures physically consistent light transport, achieving high-quality and photorealistic relighting results. Besides, the 2D-to-3D project strategy alleviates the computational cost of explicit ray tracing for visibility and global illumination, ensuring high efficiency during inference. In summary, our method makes the following key contributions:
@@
  • We present a Lighting-roubst Geometry Refinement module trained on synthetically relit multi-view data to estimate accurate depth and surface normals, effectively mitigating geometry errors caused by uneven illumination.
  • We design a Physically Grounded Neural Rendering module that combines physics-based rendering with neural rendering, achieving high-quality shading effects with improved computational efficiency.
@@ 2 Related Work
Traditional methods [debevec2000acquiring, hawkins2001photometric, debevec2002lighting, wenger2005performance, debevec2012light, weyrich2006analysis, chabert2006relighting, guo2019relightables] propose to sample the reflectance field of a human performer through a LightStage setup comprising controlled lighting systems and dense camera arrays, which can generate photorealistic renderings under novel lighting environments. However, the costly setup and complex person-specific relighting process constrain their widespread applications. With the rapid progress in neural implicit representations [mildenhall2021nerf, aliev2020neural, bi2020deep, lombardi2019neural, thies2019deferred, xu2019deep, xu2018deep], neural inverse rendering [boss2021nerd, srinivasan2021nerv, zhang2021nerfactor, zhang2021physg, boss2021neural, jin2023tensoir, yao2022neilf, zhang2023neilf++] has emerged as a promising approach for person-specific human relighting, enabling the joint recovery of geometry, material, and illumination from multi-view images captured under arbitrary lighting conditions. Recent variants [chen2022relighting4d, sun2023neural, xu2023relightable, lin2024relightable, li2024animatable, chen2024meshavatar, wang2024intrinsicavatar] extend inverse rendering to dynamic performers by integrating a parametric human body template into the learning of implicit fields, enabling temporally coherent reconstruction under varying poses and motions. However, these methods typically require time-consuming person-specific optimization and incur high computational costs, resulting in limited training and rendering speeds. With the advent of explicit point-based 3DGS [kerbl20233d] representation, recent 3DGS-based inverse rendering methods [gao2024relightable, jiang2024gaussianshader, liang2024gs, wu2025deferredgs, hong2025beam] explore the disentanglement of material, geometry, and lighting through the use of Gaussian points, enabling efficient training and fast inference. ARGS [li2024animatable] pioneers the integration of animatable 3D Gaussians and inverse rendering, delivering visually compelling and relightable full-body human avatars. RGCA [saito2024relightable] leverages dynamic relighting data captured in a LightStage and an explicit mesh template to achieve high-quality relightable head avatars. However, such methods typically rely on time-consuming person-specific optimization. Moreover, performing inverse rendering from multi-view images under arbitrary illumination is inherently ill-posed, making it particularly difficult to accurately disentangle geometry, material, and lighting, ultimately limiting the fidelity of the final relighting results.
@@
With the advancement of deep learning, data-driven methods [pandey2021total, sun2019single, zhou2019deep, wang2020single, mei2023lightpainter] have achieved impressive results in portrait relighting from a single image, primarily by leveraging convolutional neural networks trained on synthetic data derived from OLAT datasets. However, these methods typically overlook underlying human geometry, leading to physically implausible light–shadow interactions. Recent efforts [meka2020deep, ji2022geometry] attempt to address this limitation by estimating human geometry, but monocular predictions remain insufficiently accurate for reliable relighting. Other approaches [kanamori2019relighting, lagunas2021single, tajima2021relighting] attempt to learn light transport directly from 2D shading images. However, modeling complex 3D light transport in 2D space remains fundamentally challenging. Such methods are generally restricted to approximating low-frequency lighting effects and fail to generalize to unseen illumination. SwitchLight [kim2024switchlight] circumvents the need for explicit 3D geometry by introducing physics-based rendering in the image domain, achieving high-quality relighting results. Nevertheless, the absence of an explicit 3D representation limits its ability to handle occlusions and maintain spatial consistency. More recently, diffusion-based generative models [ponglertnapakorn2023difareli, ren2024relightful, zhang2025scaling] demonstrate impressive relighting performance by leveraging strong pre-trained priors and high-quality lighting datasets, producing visually striking results with high generalization ability. However, priors learned from pre-training and Light Stage supervision lack physics-based constraints during training, which in turn reduces the physical interpretability and realism of the outputs.
@@ 3 Method
As illustrated in Fig. 2, our proposed GRGS achieves generalizable and relightable human novel view synthesis via two core modules: (1) Lighting-robust Geometry Refinement (LGR) module and (2) Physically Grounded Neural Rendering (PGNR) module. The core idea is to adopt a supervised, data-driven strategy that projects intrinsic attributes from multi-view 2D observations onto 3D Gaussian representations in a feed-forward manner. Specifically, given a set of sparse-view human images, GRGS first constructs the LGR module to estimate depth using a stereo-based method and then refines per-Gaussian surface normals via a depth-aware refinement, which allows GRGS to mitigate geometry errors caused by uneven illumination (Sec. 3.1). Next, building on this accurate geometry, the PGNR module fuses physics-based rendering with neural rendering to synthesize realistic relighting phenomena. It leverages geometry-aware decoders for Gaussian parameters and intrinsic attributes, and incorporates environment map encoding to predict direct illumination scaling factors and spherical harmonic coefficients for modeling indirect illumination (Sec. 3.2). To ensure strong generalization and realistic rendering, we introduce a 2D-to-3D projection training strategy that reasonably transfers 2D illumination cues into the 3D Gaussian space, simultaneously improving rendering efficiency (Sec. 3.3).
3.1 Lighting-robust Geometry Refinement
To enable a generalizable and relightable 3D Gaussian framework, fast and accurate geometry reconstruction of the target performer is a critical foundation. This geometry not only determines the center of each Gaussian point but also serves as an essential component for the subsequent relighting stage. However, existing methods [gao2024relightable, jiang2024gaussianshader, liang2024gs, wu2025deferredgs] typically require per-scene training to optimize Gaussian centers and still struggle to guarantee high-quality and consistent geometry under various lighting. To address this challenge, we design a Lighting-roubst Geometry Refinement module, consisting of stereo-based depth estimation and depth-guided normal refinement, which jointly produce accurate geometry robust to illumination changes.
@@ 3.1 Lighting-robust Geometry Refinement
Inspired by the generalization strategy of GPS-Gaussian [zheng2024gps], we leverage RAFT-Stereo [lipson2021raft], a stereo-based depth estimator that uses disparity as a geometric constraint across views, enabling consistent depth prediction and improved generalization in diverse human-centric scenarios. However, pretrained depth estimators tend to be sensitive to illumination variations, leading to unreliable geometry estimation as shown in Fig 3. Therefore, we construct an image encoder trained by a large-scale multi-view relit dataset comprising hundreds of high-quality human scans, enabling the network to learn robust reflectance features for accurate depth estimation across diverse lighting conditions.
Specifically, given $N$ sparse-view images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ ($\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}$), captured under an arbitrary lighting of a human-centered scene, we select the two source views nearest to the target camera pose, denoted as $\{\hat{\mathbf{I}},\bar{\mathbf{I}}\}$, as input for stereo rectification [papadimitriou1996epipolar]. The rectified image pair is then passed through a shared lighting-robust feature extractor $\xi_{\text{img}}$, which is trained on the relit dataset encompassing diverse and varying illumination conditions to explicitly learn features invariant to lighting changes. This yields multi-scale radiance features $\{\hat{\mathbf{F}}_{s},\bar{\mathbf{F}}_{s}\}_{s=1}^{K}$, where $\hat{\mathbf{F}}_{s},\bar{\mathbf{F}}_{s}\in\mathbb{R}^{\frac{H}{2^{s}}\times\frac{W}{2^{s}}\times C}$ denote the feature representations at the $s$-th scale corresponding to the two input images $\{\hat{\mathbf{I}},\bar{\mathbf{I}}\}$, respectively, and $K$ is the total number of scales.
The radiance features mitigate lighting-induced feature mismatches to facilitate accurate depth estimation, as demonstrated in Sec. 4.2, while guiding the inference of geometry, material, and light transport within the 3D Gaussian representation (Sec. 3.2).
@@
$\mathbf{N}_{c}(u,v)=\frac{(\mathbf{X}(u\!+\!1,v)-\mathbf{X}(u,v))\times(\mathbf{X}(u,v\!+\!1)-\mathbf{X}(u,v))}{\|(\mathbf{X}(u\!+\!1,v)-\mathbf{X}(u,v))\times(\mathbf{X}(u,v\!+\!1)-\mathbf{X}(u,v))\|}$
where $\times$ denotes the cross product. Due to the convex upsampling, the resulting depth maps lack high-frequency geometric details and exhibit checkerboard artifacts, as illustrated in Sec. 4.2. To address this, we introduce a normal refinement module $\mathcal{N}$ based on a lightweight U-Net architecture, which leverages multi-scale radiance features $\{\mathbf{F}_{s}\}_{s=1}^{K}$ and coarse normals $\mathbf{N}_{c}$ to predict normal offsets $\Delta\mathbf{N}\in\mathbb{R}^{3}$ and recover fine-grained surface details. The refinement process can be defined as
@@
Figure 3: Geometry consistency comparison between GRGS and GPS-Gaussian under varying illumination. We visualize the normal difference heatmap, where cold colors indicate small angular deviations and hot colors denote large ones. GPS-Gaussian exhibits noticeable degradation under extreme lighting.